OpenStudy (lilai3):

WILL AWARD MEDAL/FAN!!! COMES WITH FREE OREOS AND A CERTIFICATE FOR HELPING ME!!! The product of the least and greatest of three consecutive negative odd integers is 221. Find the three integers.

OpenStudy (aayushi.somani):

-x * (-x - 4) = 221 x^2 + 4x = 221 x^2 + 4x - 221 = 0 (x + 17)(x - 13) = 0 x = -17, 13 the numbers are -13, -15, and -17.

OpenStudy (lilai3):

What are your three consecutive NEGATIVE ODD integers, if I may ask?

OpenStudy (aayushi.somani):

i wrote them -13,-15 and -17

OpenStudy (lilai3):

No, what I mean is, are your consecutive integers: -x,-x-2, and -x-4?

OpenStudy (aayushi.somani):

yeah actually..:)

OpenStudy (lilai3):

Well, one tutor (can't tell you where exactly, mind you) told me that it was -x, -x-1, and -x-3. Another told me that it was -x, -x-2, and -x-4. Which are correct? Are they both the same or not?

OpenStudy (aayushi.somani):

see u can use any..!! while u'll be getting the same answer.. !

OpenStudy (lilai3):

How did you get 13 in your answer?

OpenStudy (nikato):

no they are not the same. the first is incorrect. lets say x=5. so first integer is -5 next is -5-1 which equal -6 and -6 is not an odd integer

OpenStudy (lilai3):

So, what should it be, @nikato ?

OpenStudy (nikato):

well i was answering ur previous question, are they both the same or not? no because like i said above, -x,-x-1,-x-3 does not give us consecuive odd integers. however the other is right. -x,-x-2,-x-4 will give u consecutive odd integers

OpenStudy (lilai3):

Oh. I see. So what @aayushi.somani was right?

OpenStudy (nikato):

correct. and to answer ur other question (x + 17)(x - 13) = 0 which was in his/her work and u can seperate to x+17=0 or x-13=0 so x=-17 or x=13 but since the question asks for ODD integer, u can disgard the positive 13

OpenStudy (nikato):

sorry, i accidentally closed this window and had to reopen it

OpenStudy (lilai3):

Oh, it's alright. LOL. As long as you are helping me, obviously, it's alright. xD

OpenStudy (nikato):

ok.. no problem

OpenStudy (lilai3):

So even if your answer is positive 13, then you could just take it for a negative 13?

OpenStudy (aayushi.somani):

no.. the positive 13 has be rejected.. and we have the keast negative value -17 soo v will now count for 2 more consecutive odd negative no.

OpenStudy (lilai3):

Okay. Thank you